Help

I. Over the past 300 days, DiCarlo has experienced 54 days with no automobiles sold, 117 days with 1 automobile sold, 72 days with 2 automobiles sold, 42 days with 3 automobiles sold, 12 days with 4 automobiles sold, and 3 days with 5 automobiles sold. Suppose we consider the experiment of observing a day of operations at DiCarlo Motors and define the random variable of interest as x the number of automobiles sold during a day. a) Use the relative frequency method to develop a probability distribution for the number of cars sold per day at DiCarlo Motors in Saratoga, New York. b) The expected value for the number of automobiles sold during a day c) The variance for the number of automobiles sold during a day
Answer #1

Solution

Given x = number of automobiles sold during a day, the given frequency and the related requisite calculations are shown in the following table.

i

Number of automobiles sold during the day

(xi)

Number of days

(Frequency)

(fi)

Relative

Frequency

Ri = p(xi)

= fi/300

xi.p(xi)

xi2.p(xi)

1

0

54

0.18

0.00

0.00

2

1

117

0.39

0.39

0.39

3

2

72

0.24

0.48

0.96

4

3

42

0.14

0.42

1.26

5

4

12

0.04

0.16

0.64

6

5

3

0.01

0.05

0.25

Total

300

1.00

1.50

3.50

Part (a)

Probability Distribution

Number of automobiles sold during the day

(xi)

Probability

0

0.18

1

0.39

2

0.24

3

0.14

4

0.04

5

0.01

Total

1.00

ANSWER

Part (b)

Expected number of cars sold per day = E(X) = ∑[i = 1, 6]{xi.p(xi)} = 1.50 ANSWER [refer to column 5 of table at the top]

Part (c)

Variance of the number of cars sold per day = V(X)

= E(X2) – { E(X)}2

= ∑[i = 1, 6]{xi2.p(xi)} – 1.52 [refer to the answer of Part (b)]

= 3.5 – 2.25 [refer to column 6 of table at the top]

= 1.25 ANSWER